Prev Next

SW&IT Meeting Minutes 2020-05-18

Attendees: Carl, David, Igor, Matthieu, Stephen, Xianyu, Yang


Labor Reporting

Report auto-generated every Monday at 6:30am, Do not edit this table, Update your entries in smartsheets.

LR20200504 - Bec, Matthieu 80.0
LR20200504 - Holmberg, Carl 40.0
LR20200504 - Edmeades, David 40.0
LR20200504 - Sola, Igor 40.0
LR20200504 - Hooper, Stephen 80.0
LR20200504 - Zhang, Yang 38.0
LR20200504 - Zhang, Xianyu 80.0

News from management

  • student Ad out for Iraf to Python3 conversion
  • Senior Ad WIP (into Gene's hands)

Previous Action Items

Owner Action Due Date Completion Date

Open topics

Team reports


  • ROWG monitor, room 507 work and stuck Gnome on new Fedora 31

  • DNS in Tucson
    • moved AD to
    • unbound at
    • issues with IP stack on virtio-net(?) and aliases
    • changed AD for VPN, rm507-2 and developer machines
  • moved developer machines to vm1.tucson: isola-luci, mbec-dev, xzhang-soul-adsec, yzhang-dev
  • Synology issues
  • meetings


  • mountain power on CR
  • laptops for Leroy
  • play with networking on ovirt and VMWare
    • get dns_recurser to listen to
    • get dns_recurser to allow for NIC
  • authoritative DNS for Tucson
  • satelite images for info


  • Jim printer purchase
  • Jim O365... thing? I think his auth token had expried and he just needed to reauthenticate via shibboleth
  • Reopened idea of upgrading Luci servers
    • Have heard that science views this in a positive light
    • My only objection was based in monetary ignorance
  • Start cataloging data usage on Synology
  • Hearing rumblings about dodgy home internet connections; any action on our part?
  • Cisco training purchased
    • Slight hiccup; Louise is fixing
  • Crucial SSD RMA adventure
    • I am assured that a second replacement is on its way to central receiving
  • On hold re: laser cutter computer
  • Synology still hiccupping; SH said he's working on it
  • GR request a look into adding more robust search options to our IssueTrak install
  • Work with IS to refine new luci computer spec; took advantage of today's computer restart to answer the questions we'd built up.


  • Previous:
    • AO-Cravat refactoring and debugging (some possible progress based on some of Al's feedback, will meet with Al/Steve end of the Month).
    • TMS information gathering (produce a sketch and organize a meeting).
    • modsAlign refactoring
  • Next:
    • AO-Cravat refactoring/debugging continues (organize a zoom meeting/Doodle poll).
    • Finish TMS sketch and possibly organize a zoom meeting.
    • TCS/IIF VCAN doc issue.
    • Some modsAlign refactoring if have time.
    • Some MODS WCS editor work when have time.
    • Catch up with safety training.


  • Last week
    • Merge UAO to SOUL
    • test the merged AO versions on the VM simulator
    • PB SW/HW discussion
    • Review SOUL PDR document
    • Meetings
  • This week
    • troubleshooting of SX ASM issue
    • working on PB action items
    • Merge UAO to SOUL


  • Last week
    • OC, Management
    • Technical Requirements for a Remote Observing Facility - draft v0.1
    • DMS
      • Satellite web component, Weather loss explorer dashboard
      • misc. refactoring
      • tiemstamp issue with allsky and pre-2019 directory structure
    • AO PB/SOUL/ARGOS/LN, draft presentations v0.2
    • Effective use of IT (Bucki's training)
  • This week
    • ROWG #7 meeting
    • SHARK-VIS follow up new documents, send new IIF document and clarify DD entries for SOUL
    • DMS
      • websockets, AdSec, etc.
      • LBC stiched png preview
    • cobbler / kickstart, ims f30


Last Week:
  • DMS: complete imonnit ingest changes.
  • OA Control: Research software changes needed for encoder support binary output.
  • Safety training
This Week:
  • OA Control: Install DT dev IDE, document existing PMAC code.
  • Safety training


Last week:
  • LUCI
    • Worked a bit more in the LUCI environment in the virtualbox vm Centos7. I can now use it locally.
  • OT
    • Went through a couple of bugs Barry reported, still need to understand the MODS twilights flats script generation.
This week:
  • LUCI
    • support Dave and James in the LUCI1 detector testing.
    • work on #IT7952 an RTD/Aladin GUI issue.
    • Look into LUCI WebIO device configuration.
  • PIT
    • If I find some time, work on the PIT admin tool backend.
  • OT
    • We are aiming at a release possibly this week?
Topic revision: r8 - 18 May 2020, DavidEdmeades
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback